OpenStudy (anonymous):

if you are given a^2 ,b^2 and value of k .you are asked to find out c^2 not using calculator or multiply cxc=c^2 if three numbers are ;a=a ,b=a+k ,c=b+k;here k is a constant.

OpenStudy (anonymous):

then work backwards

OpenStudy (anonymous):

\[c^2= (b+k)^2\]

OpenStudy (anonymous):

\[(b+k)^2 = b^2 +2bk +k^2\]

OpenStudy (anonymous):

Actually I discover a equation rule :rule 1.c^2=2xb^2-a^2+k^2 rule 2; c^2=2xcxk+b^2-k^2 rule 3; c^2 =(2xb -a)^2 . your answer is right ,you just simply say the answer . I ask you this question to know alternative rule to find out c^2 . Please justify my rule 1,2,3. thanks open study .if my rule is proved please sent me an email at raselazam23@gmail.com. Thanks open study.